OpenStudy (agentnao):

I know this is a chemistry question, but since there is no chemistry group, I'm going to ask it here. A 230.0-mL sample of a 0.275 M solution is left on a hot plate overnight; the following morning the solution is 1.15 M. What volume of solvent has evaporated from the 0.275 M solution?

OpenStudy (anonymous):

there is a chemistry section.

OpenStudy (dumbcow):

what is the solution consist of? different elements have different mole to volume ratios i believe

OpenStudy (anonymous):

but, \[M _{1}V _{1}=M _{2}V _{2}\] so (230)(.275) = (1.15)(x) x = 63.25/(1.15) x = 55 mL

OpenStudy (agentnao):

There might be, but I didn't see it. and it doesn't say dumbcow. I don't think it's necessary for this problem.

OpenStudy (agentnao):

Whoa. Holy cow. That was really quick and simple iamtheos! Thank you! That was so simple actually, that I feel like I should have known it..

OpenStudy (agentnao):

But since I didn't, thank you VERY MUCH.

OpenStudy (anonymous):

you're welcome :) and to find the Chemistry section, go to the home page and click "View All Groups". Chemistry is on the left side.

OpenStudy (agentnao):

Ohh, I see. Alright, cool! But since you were so helpful with this question, can I ask you one more?

OpenStudy (anonymous):

Of course.

OpenStudy (agentnao):

What mass of calcium chloride, CaCl2, is needed to prepare 2.000 L of a 1.37 M solution?

OpenStudy (agentnao):

Is this also a very simple solution to this as well??

OpenStudy (agentnao):

*there

OpenStudy (anonymous):

Molarity = M and M's units are mol/L so, 1.37 = mol/2 mol of CaCl2 = 2.74 then use the molar mass of Calcium Chlorate (101g) so (101)(2.74) = 276.74 Not as easy, but still not too bad :D

OpenStudy (agentnao):

Hm...but that's not one of the options on the multiple choice. My friend got 304 and I got something totally wrong. (304 IS one of the options though)

OpenStudy (anonymous):

check your numbers, did you copy it to the question correctly?

OpenStudy (agentnao):

Yep!

OpenStudy (agentnao):

I copy and pasted it from the direct source.